# Neural Signals Generate Clinical Notes in the Wild

### v1

@@ Abstract
Generating clinical reports that summarize abnormal patterns, diagnostic findings, and clinical interpretations from long-term EEG recordings remains labor-intensive. We curate a large-scale clinical EEG dataset with 9,9229{,}922 reports paired with approximately 11,00011{,}000 hours of EEG recordings from 9,0489{,}048 patients. We therefore develop CELM, the first clinical EEG-to-Language foundation model capable of summarizing long-duration, variable-length EEG recordings and performing end-to-end clinical report generation at multiple scales, including recording description, background activity, epileptiform abnormalities, events/seizures, and impressions. Experimental results show that, with patient history supervision, our method achieves $70\%$$95\%$ average relative improvements in standard generation metrics (e.g., ROUGE-1 and METEOR) from 0.20.20.30.3 to 0.40.40.60.6. In the zero-shot setting without patient history, CELM attains generation scores in the range of 0.430.430.520.52, compared to baselines of 0.170.170.260.26. CELM integrates pretrained EEG foundation models with language models to enable scalable multimodal learning. We release our model and benchmark construction pipeline at
@@ 1 Introduction
To support this profound clinical practice, several automated clinical EEG report generation models based on advanced deep learning techniques have been proposed (Biswal et al., 2019; Lee et al., 2023; Yin & Shin, 2025). These methods aim to encode EEG recordings into latent representations that capture temporal structure and clinically relevant phenotypes, which are then decoded into clinical reports. However, existing methods remain limited in several aspects. ① Conceptually, some works formulate report generation as a phenotype classification or retrieval problem followed by text decoding. They do not model EEG-to-report generation as an end-to-end learning task, and the misalignment between classification and generation objectives does not guarantee optimal report generation. ② Methodologically, these methods operate on short EEG segments and fixed-context templates. They fail to model long-term temporal context and global diagnostic reasoning required for interpreting long-duration EEG data. Moreover, fixed-context templates prevent multi-granularity report generation beyond predefined categories. ③ Practically, these methods are developed as task-specific models, each tailored to a narrow reporting objective (e.g., impression-only reports) (Lee et al., 2023). In contrast, clinical practice requires multi-level reporting, where neurologists routinely construct overall EEG summaries and section-wise outputs, including impressions, interpretations, and event or seizure annotations.
@@
Our motivation stems from recent advances in multimodal foundation models that enable text generation from non-text modalities (e.g., images and speech) (Bordes et al., 2024; Shen et al., 2025). Recent studies have begun to incorporate EEG signals into language-based modeling frameworks (Ndir et al., 2025; Gijsen & Ritter, 2025). However, they primarily leverage text supervision to enhance EEG classification and do not support EEG-to-text report generation.
@@
Experimental results show that CELM consistently outperforms all baselines across all settings. With patient history, it achieves $70\%$$95\%$ average relative improvements, with standard generation metrics (e.g., METEOR (Banerjee & Lavie, 2005) and ROUGE-1 (Lin, 2004)) from 0.20.20.30.3 to 0.40.40.60.6 across diverse cases. In the zero-context setting without patient history, CELM achieves generation scores in the range of 0.430.430.520.52, compared to baselines of 0.170.170.260.26. We provide multi-perspective case analyses to shed light on future ELM development.
@@ 2 Related Work
Existing work at the intersection of EEG and natural language can be broadly categorized into two paradigms: (1) EEG-to-language decoding and (2) text-enhanced EEG representation learning. The goal of EEG-to-language decoding is to reconstruct textual content from concurrent EEG recordings of subjects reading or imagining speech (Herff et al., 2015; Wang & Ji, 2022). This line of work spans both invasive approaches using electrocorticography (ECoG) (Makin et al., 2020; Willett et al., 2023) and non-invasive methods based on scalp EEG (Wang & Ji, 2022; Duan et al., 2023; Zhou et al., 2024). However, these methods assume precise EEG–text alignment, whereas clinical EEG consists of heterogeneous events embedded in long, continuous recordings spanning hours to days (Sun et al., 2025; Zafar et al., 2025). The second paradigm leverages clinical notes as auxiliary supervision to enhance EEG representation learning, rather than decoding language directly from neural activity. Inspired by vision–language pretraining frameworks (Radford et al., 2021), a recent work  (Ndir et al., 2025) aligns EEG data with clinical report text in a shared feature space. However, it remains focused on discriminative objectives and does not support automated clinical report generation.
Clinical EEG Report Generation. An early attempt at EEG report generation, EEGtoText (Biswal et al., 2019), proposed a two-stage pipeline that first classifies EEG phenotypes, and then generates report text via a text decoder conditioned on the classified labels. However, this relies on intermediate phenotyping as a bottleneck, limiting the capabilities to capture nuanced clinical findings beyond predefined categories. While some attempts jointly learn EEG encoders and text decoders (Lee et al., 2023), they rely on fixed segmentation and template-based generation (Yin & Shin, 2025). As a result, these methods fall short of end-to-end clinical report generation from long-duration EEG recordings. We address the above challenges by introducing the first family of EEG–Language foundation models.
@@ 5 Experiments and Results
As no existing EEG multimodal models with generative capabilities are available for EEG report generation, we establish baselines by augmenting local LLMs with EEG-derived inputs. We evaluate two settings: (1) Unimodal + text-only input, where the LLMs receive only clinical context as input without EEG data, providing a lower bound that quantifies hallucination and measures performance when the model does not attend to neural signals, and (2) Unimodal + text and EEG features input, where we augment the clinical context with channel-wise spectral band power features (delta, theta, alpha, beta, gamma) extracted from each EEG session, addressing context length limitation that prevents providing raw EEG arrays directly. We evaluate LLMs from three model families, including LLaMA-3 (Grattafiori et al., 2024), Gemma-3 (Team et al., 2025), and Qwen-3 (Yang et al., 2025), across multiple parameter scales. Additionally, we include MedGemma (Sellergren et al., 2025) as a medical LLM baseline. Additional implementation details are provided in Appendix B. We report smoothed BLEU-1 (Papineni et al., 2002), ROUGE-1 (Lin, 2004), and METEOR (Banerjee & Lavie, 2005) scores to evaluate report generation performance. Comprehensive results and metrics are provided in Appendix D.

### v2

@@ Abstract
Generating clinical reports that summarize abnormal patterns, diagnostic findings, and clinical interpretations from long-term EEG recordings remains labor-intensive. We curate a large-scale clinical EEG dataset with 9,9229{,}922 reports paired with approximately 11,00011{,}000 hours of EEG recordings from 9,0489{,}048 patients. We therefore develop CELM, the first clinical EEG-to-Language foundation model capable of summarizing long-duration, variable-length EEG recordings and performing end-to-end clinical report generation at multiple scales, including recording description, background activity, epileptiform abnormalities, events/seizures, and impressions. Experimental results show that, with patient history supervision, our method achieves $70\%$$95\%$ average relative improvements in standard generation metrics (e.g., ROUGE-1 and METEOR) from 0.20.20.30.3 to 0.40.40.60.6. In the zero-shot setting without patient history, CELM attains generation scores in the range of 0.430.430.520.52, compared to baselines of 0.170.170.260.26. CELM integrates pretrained EEG foundation models with language models to enable scalable multimodal learning. We release our model and benchmark construction pipeline at https://github.com/Jathurshan0330/CELM.
@@ 1 Introduction
To support this profound clinical practice, several automated clinical EEG report generation models based on advanced deep learning techniques have been proposed (Biswal et al., 2019; Lee et al., 2023; Yin and Shin, 2025). These methods aim to encode EEG recordings into latent representations that capture temporal structure and clinically relevant phenotypes, which are then decoded into clinical reports. However, existing methods remain limited in several aspects. ① Conceptually, some works formulate report generation as a phenotype classification or retrieval problem followed by text decoding. They do not model EEG-to-report generation as an end-to-end learning task, and the misalignment between classification and generation objectives does not guarantee optimal report generation. ② Methodologically, these methods operate on short EEG segments and fixed-context templates. They fail to model long-term temporal context and global diagnostic reasoning required for interpreting long-duration EEG data. Moreover, fixed-context templates prevent multi-granularity report generation beyond predefined categories. ③ Practically, these methods are developed as task-specific models, each tailored to a narrow reporting objective (e.g., impression-only reports) (Lee et al., 2023). In contrast, clinical practice requires multi-level reporting, where neurologists routinely construct overall EEG summaries and section-wise outputs, including impressions, interpretations, and event or seizure annotations.
@@
Our motivation stems from recent advances in multimodal foundation models that enable text generation from non-text modalities (e.g., images and speech) (Bordes et al., 2024; Shen et al., 2025). Recent studies have begun to incorporate EEG signals into language-based modeling frameworks (Ndir et al., 2025; Gijsen and Ritter, 2025). However, they primarily leverage text supervision to enhance EEG classification and do not support EEG-to-text report generation.
@@
Experimental results show that CELM consistently outperforms all baselines across all settings. With patient history, it achieves $70\%$$95\%$ average relative improvements, with standard generation metrics (e.g., METEOR (Banerjee and Lavie, 2005) and ROUGE-1 (Lin, 2004)) from 0.20.20.30.3 to 0.40.40.60.6 across diverse cases. In the zero-context setting without patient history, CELM achieves generation scores in the range of 0.430.430.520.52, compared to baselines of 0.170.170.260.26. We provide multi-perspective case analyses to shed light on future ELM development.
@@ 2 Related Work
Existing work at the intersection of EEG and natural language can be broadly categorized into two paradigms: (1) EEG-to-language decoding and (2) text-enhanced EEG representation learning. The goal of EEG-to-language decoding is to reconstruct textual content from concurrent EEG recordings of subjects reading or imagining speech (Herff et al., 2015; Wang and Ji, 2022). This line of work spans both invasive approaches using electrocorticography (ECoG) (Makin et al., 2020; Willett et al., 2023) and non-invasive methods based on scalp EEG (Wang and Ji, 2022; Duan et al., 2023; Zhou et al., 2024). However, these methods assume precise EEG–text alignment, whereas clinical EEG consists of heterogeneous events embedded in long, continuous recordings spanning hours to days (Sun et al., 2025; Zafar et al., 2025). The second paradigm leverages clinical notes as auxiliary supervision to enhance EEG representation learning, rather than decoding language directly from neural activity. Inspired by vision–language pretraining frameworks (Radford et al., 2021), a recent work  (Ndir et al., 2025) aligns EEG data with clinical report text in a shared feature space. However, it remains focused on discriminative objectives and does not support automated clinical report generation.
Clinical EEG Report Generation. An early attempt at EEG report generation, EEGtoText (Biswal et al., 2019), proposed a two-stage pipeline that first classifies EEG phenotypes, and then generates report text via a text decoder conditioned on the classified labels. However, this relies on intermediate phenotyping as a bottleneck, limiting the capabilities to capture nuanced clinical findings beyond predefined categories. While some attempts jointly learn EEG encoders and text decoders (Lee et al., 2023), they rely on fixed segmentation and template-based generation (Yin and Shin, 2025). As a result, these methods fall short of end-to-end clinical report generation from long-duration EEG recordings. We address the above challenges by introducing the first family of EEG–Language foundation models.
@@ 5 Experiments and Results
As no existing EEG multimodal models with generative capabilities are available for EEG report generation, we establish baselines by augmenting local LLMs with EEG-derived inputs. We evaluate two settings: (1) Unimodal + text-only input, where the LLMs receive only clinical context as input without EEG data, providing a lower bound that quantifies hallucination and measures performance when the model does not attend to neural signals, and (2) Unimodal + text and EEG features input, where we augment the clinical context with channel-wise spectral band power features (delta, theta, alpha, beta, gamma) extracted from each EEG session, addressing context length limitation that prevents providing raw EEG arrays directly. We evaluate LLMs from three model families, including LLaMA-3 (Grattafiori et al., 2024), Gemma-3 (Team et al., 2025), and Qwen-3 (Yang et al., 2025), across multiple parameter scales. Additionally, we include MedGemma (Sellergren et al., 2025) as a medical LLM baseline. Additional implementation details are provided in Appendix B. We report smoothed BLEU-1 (Papineni et al., 2002), ROUGE-1 (Lin, 2004), and METEOR (Banerjee and Lavie, 2005) scores to evaluate report generation performance. Comprehensive results and metrics are provided in Appendix D.